OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. (9 − x)e^(x) dx

OpenStudy (raden):

u = 9 - x du = - dx or dv = e^x dx v = int e^x dx = e^x now apply into formula int u dv = uv - int vdu